### v1

@@ Abstract
In the area of computer assisted IVF-ET, existing researches mainly focus on the morphological grading of embryos. As shown in Fig. 1, we can obtain microscope images of embryonic development between step 3 and step 4 of IVF-ET. In order to perform the embryos morphological grading task, [Wu et al.(2021)Wu, Yan, Li, Li, Wang, Chang, Yu, Jin, Ma, Luo, et al.] and [Liu et al.(2023a)Liu, Li, Dai, Shan, Zhang, Zhuang, Lee, Wong, Yue, Huang, et al.] apply convolutional neural networks (CNN) based on static embryo images; [Wang et al.(2024)Wang, Wang, Gao, Chen, Gao, Ma, Jiang, Yang, Feng, Zhang, et al.] and [Lukyanenko et al.(2021)Lukyanenko, Jang, Wei, Struyven, Kim, Leahy, Yang, Rush, Ben-Yosef, Needleman, et al.] apply transformer and two-stream neural network based on time-lapse microscopy (TLM) images, respectively. In addition, [Cheng et al.(2024)Cheng, Chen, Hu, Meng, Liu, Chen, Wu, and Ying] fuse multi-focal images to predict outcome of blastocyst. The performance of these methods outperform laboratory technicians, because embryos grading is completely based on morphological information and the salient morphological characteristics among different grades are easily distinguishable for machines. However, morphological grading is indirect and less relevant to the pregnancy outcome of IVF-ET as discussed in the previous paragraph. Therefore, more researches tend to predict the pregnancy outcome directly.
@@ 2 Method
In the field of medical multi-modal fusion, the final fusion features are obtained by complementing the unique features and enhancing the common features of different modalities. However, there is a complex relationship between the features of the same modality and different modalities, which is not a simple linear relationship. So it is difficult to be captured by the model. Inspired by the decoupling operation in [Dong et al.(2023)Dong, Nejjar, Sun, Chatzi, and Fink] [Li et al.(2023a)Li, Fei, Liao, Zhao, Teng, seng Chua, Ji, and Li], we use the decoupling fusion strategy explicitly decouples the features of different modalities into unique and common features, transforming modal features from disorder to order, simplifying the relationships between features and better modeling the complex interactions between modalities. So we propose the decoupling fusion module, a simple and effective multi-modal fusion module as showed in Fig. 2. The input information of the model are temporal grayscale embryo images and fertility table indicator. Embryo images are denoted as , where denote different days, and and denote the height and the width of an image, respectively. Table indicators are denoted as , where denotes the number of indicators.
@@ 2.1 Image Extractor
To fuse the first three days embryo images for pregnancy prediction, we design a temporal image fusion network (image extractor). This network consists of three parts as shown in Fig. 3: a backbone aiming at extracting image features; a spatial-temporal position encoding (STPE) to obtain spatial information within a single image and temporal information among different images; and a Transformer [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] that combine STPE to fuse temporal image features. Firstly, we use the backbone to extract embryo image features from the first three days. Then, we reshape the image features into tokens and add positional information to them. We encode image features as spatial position encoding through a convolution layer: . By using convolution operation to capture local spatial features, we can obtain spatial position information within a single image [Chu et al.(2021)Chu, Tian, Zhang, Wang, and Shen]. Image features are encoded as temporal position encoding through a pooling layer and a series of recurrent layers: . By using recurrent operations to capture temporal dependencies, we can obtain temporal position information among different images [Hochreiter and Schmidhuber(1997)]. To align image features, the needs to be replicated to get . The and are integrated through a position encoding attention:
@@ 2.3 Decoupling Fusion Module
where is the L1-norm, is the dimensionality of the feature. We obtain the final pregnancy prediction result by fusing the decoupled common and unique features: . Finally, we apply cross entropy loss to minimize the difference between the predicted results and the true labels. The classification loss is as follows:
@@ Appendix C Dataset Samples
In the Fig.A.2 and Table A.1, we present some examples of image modality and table modality used for pregnancy prediction, respectively. The image is obtained by cropping the excess blank space in the original image. We replace the missing value in the table data with the average value and normalize the tabular features with min-max before inputting them into the model.
@@ Appendix E Interpretability
In order to analyze the reasons for the success of the model, we conduct interpretability analysis on the model, mainly manifest in two aspects. Firstly, as shown in Fig. A.4, we use SHAP [Lundberg and Lee(2017)] to output the importance ranking of clinical indicators in our model, which focuses more on features such as female age, high-quality embryos and so on. Among them, A.4(a) ranks the importance of table features. A.4(b) is a beeswarm, which depicts the SHAP values of each sample under different features. Secondly, we use the Grad-Cam [Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra] to visualize the class activation maps of the first three days of embryonic development images in Fig. A.5. We can see the areas that our model focuses on are the edges of embryonic cells. These interpretable results are consistent with the experience of obstetricians and gynecologists.

### v2

@@ Abstract
In the area of computer assisted IVF-ET, existing researches mainly focus on the morphological grading of embryos. As shown in Fig. 1, we can obtain microscope images of embryonic development between step 3 and step 4 of IVF-ET. In order to perform the embryos morphological grading task, [Wu et al.(2021)Wu, Yan, Li, Li, Wang, Chang, Yu, Jin, Ma, Luo, et al.] and [Liu et al.(2023a)Liu, Li, Dai, Shan, Zhang, Zhuang, Lee, Wong, Yue, Huang, et al.] apply convolutional neural networks (CNN) based on static embryo images; [Wang et al.(2024)Wang, Wang, Gao, Chen, Gao, Ma, Jiang, Yang, Feng, Zhang, et al.] and [Lukyanenko et al.(2021)Lukyanenko, Jang, Wei, Struyven, Kim, Leahy, Yang, Rush, Ben-Yosef, Needleman, et al.] apply transformer and two-stream neural network based on time-lapse microscopy (TLM) images, respectively. In addition, [Cheng et al.(2024)Cheng, Chen, Hu, Meng, Liu, Chen, Wu, and Ying] fuse multi-focal images to predict grade of blastocyst. The performance of these methods outperform laboratory technicians, because embryos grading is completely based on morphological information and the salient morphological characteristics among different grades are easily distinguishable for machines. However, morphological grading is indirect and less relevant to the pregnancy outcome of IVF-ET as discussed in the previous paragraph. Therefore, more researches tend to predict the pregnancy outcome directly.
@@ 2 Method
In the field of medical multi-modal fusion, the final fusion features are obtained by complementing the unique features and enhancing the common features of different modalities. However, there is a complex relationship between the features of the same modality and different modalities, which is not a simple linear relationship. So it is difficult to be captured by the model. Inspired by the decoupling operation in [Dong et al.(2023)Dong, Nejjar, Sun, Chatzi, and Fink] [Li et al.(2023a)Li, Fei, Liao, Zhao, Teng, seng Chua, Ji, and Li], we use the decoupling fusion strategy explicitly decouples the features of different modalities into unique and common features, which is a shift from entangled features to disentangled ones, simplifying the relationships between features and better modeling the complex interactions between modalities. So we propose the decoupling fusion module, a simple and effective multi-modal fusion module as showed in Fig. 2. The input information of the model are temporal grayscale embryo images and fertility table indicator. Embryo images are denoted as , where denote different days, and and denote the height and the width of an image, respectively. Table indicators are denoted as , where denotes the number of indicators.
@@ 2.1 Image Extractor
To fuse the first three days embryo images for pregnancy prediction, we design a temporal image fusion network (image extractor). This network consists of three parts as shown in Fig. 3: a backbone aiming at extracting image features; a spatial-temporal position encoding (STPE) to obtain spatial information within a single image and temporal information among different images; and a Transformer [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] that combine STPE to fuse temporal image features. Firstly, we use the backbone to extract embryo image features from the first three days, where is the number of channels and is the scaling factor. Then, we reshape the image features into tokens and add positional information to them. We encode image features as spatial position encoding through a convolution layer: . By using convolution operation to capture local spatial features, we can obtain spatial position information within a single image [Chu et al.(2021)Chu, Tian, Zhang, Wang, and Shen]. Image features are encoded as temporal position encoding through a pooling layer and a series of recurrent layers: . By using recurrent operations to capture temporal dependencies, we can obtain temporal position information among different images [Hochreiter and Schmidhuber(1997)]. To align image features, the needs to be replicated to get . The and are integrated through a position encoding attention:
@@ 2.3 Decoupling Fusion Module
where is the L1-norm, is the dimensionality of the feature, and are decoder. We obtain the final pregnancy prediction result by fusing the decoupled common and unique features: , which is composed of three layers of MLP. Finally, we apply cross entropy loss to minimize the difference between the predicted results and the true labels. The classification loss is as follows:
@@ Appendix C Dataset Samples
In the Fig.A.2 and Table A.1, we present some examples of image modality and table modality used for pregnancy prediction, respectively. When processing image data, we use the following data enhancement during training and testing: resize to 256 pixels and then center cropping to 224 pixels. Finally, normalization with a mean of 0.566 and a variance of 0.063 is used, and the mean and variance were obtained by statistics of the whole data set. When processing tabular data, we use the average of features to replace missing features. Then we input the features into the neural network after min-max normalization.
@@ D.4 Uni-modal Experiment on The ODIR
D.5 Generalization of Other Dataset
We collect a dataset of 218 cases from Guangzhou Women and Children’s Medical Center as an additional test set to test our DeFusion model, with 56 pregnant cases and 162 non-pregnant cases in this dataset. As shown in the Table A.5, our model has certain generalization ability without fine tuning.
| Modality | Method | AUC | F1 | Accuracy |
| Image+Table | Ours(DeFusion) | 0.616 | 0.660 | 0.642 |
@@ Appendix E Interpretability
In order to analyze the reasons for the success of the model, we conduct interpretability analysis on the model, mainly manifest in two aspects. Firstly, as shown in Fig. A.4, we use SHAP [Lundberg and Lee(2017)] to output the importance ranking of clinical indicators in our model, which focuses more on features such as female age, high-quality embryos and so on. Among them, A.4(a) ranks the importance of table features. A.4(b) is a beeswarm, which depicts the SHAP values of each sample under different features.
Secondly, we use the Grad-Cam [Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra] to visualize the class activation maps of the first three days of embryonic development images in Fig. A.5. We can see the areas that our model focuses on are the edges of embryonic cells. These interpretable results are consistent with the experience of obstetricians and gynecologists.
Appendix F Limitation
Although our study is superior to other methods, there are limitations to our method and data set. For our method, our premise is that there is correlation between the multi-modal data. If there is no correlation between the multi-modal data, our method may not work well. In addition, when decoupling unique features and common features, we only use a simple cross-reconstruction loss constraint, which is very weak. Although the decoupling visualization by t-SNE and Pearson correlation coefficient proves the effectiveness of the decoupling method, adding stronger loss function constraints to the decoupling process may make the decoupling process smoother. Finally, because we add the decoupling module to the multi-modal fusion process, our computational complexity is higher than the simple ADD fusion, which is not conducive to our deployment of the model to the end-to-end device.
For our dataset, we only collected 4046 cases of data, which is not enough for deep learning. And, there are some missing values in our clinical indicator data, which is unfavorable to the prediction of results. In addition, our image data is three images taken every other day using a normal microscope, and many studies now use time-lapse microscopes, which can acquire images at the hour or even minute level, so our image data lacks a lot of temporal information compared to other studies.